OpenStudy (cenaida):

Solve the following system of equations by substitution. x + y = 5 y = 3 A. (-2, 3) B. (2, -3) C. (0, 0) D. (2, 3)

OpenStudy (luigi0210):

Sub in y x+3=5 Now solve for x:

OpenStudy (cenaida):

2?

OpenStudy (luigi0210):

Yup, so (2, y) and y is given as: 3 So...

OpenStudy (cenaida):

so the answer is D! aww thanx luigy! <3 im gonna miss u so much ur the only one that i understand when u explain to me :) i love u pal!

OpenStudy (luigi0210):

No problem, and there's always nin ;)

OpenStudy (nincompoop):

May I jump right in? it seems intuitive that we can tell 2 is the right answer, but we want to make sure that we show systematically how that is so. x+y = 5; y =3 x+3=5 x+3-3=5-3 x+0 = 2 x=2;y3 (2,3)

OpenStudy (luigi0210):

^ See

OpenStudy (cenaida):

hahaha oh i love you guys... :)